# Chain-of-Thought Reasoning Improves Context-Aware Translation with Large Language Models

### v1

@@ 7.  Conclusion
This paper evaluated the capacity of several LLMs to pass two benchmarks for contextual MT, one targeting pronoun translation and the other one lexical disambiguation and coherence. We compared several types of prompts, and showed that CoT prompting elicits reasoning that leads to the best results. We found that the best models improve considerably the state of the art on the contrastive task, with accuracies slightly above 0.95 for both targeted phenomena. They also score highly on the four translation quality metrics, with COMET scores of 0.92 and 0.89. Moreover, we observed a “wise get wiser” effect, as the improvement brought by reasoning is positively correlated to the scores of the same LLMs without reasoning. In other words, the strongest models are also those that benefit the most from reasoning.
@@ 8.  Bibliographical References
[…]
@@ A.1.  Prompt Variants
This section presents the full text of the prompts used in our contrastive evaluation. Each configuration has a system prompt (except when ‘none’ is indicated) and a user prompt. The various prompts differ in their structures and instruction levels, while maintaining a consistent task format. Placeholders appear in curly braces in the prompt templates: {context_en}, {source_sentence}, {context_fr}, {option1}, {option2}. In the contrastive setting, two candidate translations are presented, and the model must output a constrained choice, either as “Choice: 1” or as “Choice: 2”.
(none, simple)
System prompt
@@
(simple, simple)
@@
(detailed, simple)
@@
(simple, step-by-step) - Anaphora
@@
(simple, step-by-step) - Lexical Choice
@@ A.2.  Results: Anaphora
A.2.1.  Accuracy and Inconsistency Scores, Elapsed Time and Cost
This section reports results on the contrastive anaphora benchmark, for each LLM and prompt configuration.
Model (none, simple) (simple, simple) (detailed, simple) (simple, step-by-step)
@@ A.2.1.  Accuracy and Inconsistency Scores, Elapsed Time and Cost
Table 1: Accuracy (acc $\uparrow$) and inconsistency (inc $\downarrow$) across models and prompts on the anaphora validation set (separated by ‘/’ in each cell). The systems are ranked by decreasing accuracy of the best overall prompt, which is the (simple, step-by-step) prompt.
Model (none, simple) (simple, simple) (detailed, simple) (simple, step-by-step)
@@
Table 2: Mean elapsed time per prompt (seconds) across prompt configurations on the anaphora validation set. Open-source (Ollama) models listed first, followed by OpenAI models.
Model (none, simple) (simple, simple) (detailed, simple) (simple, step-by-step)
@@
Table 3: Total API cost (USD) by prompt configuration for OpenAI models on the anaphora validation set. Experiments were run in spring 2025; reported costs reflect that period’s pricing and our token usage.
@@
Table 4: Comparison of accuracy (Acc $\uparrow$) and inconsistency (Inc $\downarrow$) on validation and test sets for the anaphora task, across prompt configurations. Validation uses the first 100 couples and reproduces the scores from Table 1 above; test uses the held-out 100 couples. Each couple is presented in both option orders (correct option first/second), and scores are averaged across the two presentations.
A.2.2.  Graphical Representation of Scores
Figure 1: Mean accuracy (top) and inconsistency (bottom; lower is better) across prompt configurations for each model on the anaphora validation set. Reproduced here from the main text to keep the appendix self-contained.

### v2

@@ 7.  Conclusion
This paper evaluated the capacity of several LLMs to pass two benchmarks for contextual MT, one targeting pronoun translation and the other one targeting lexical coherence. We compared several types of prompts, and showed that CoT prompting elicits reasoning steps that lead to the best results. We found that the best models improve considerably the state of the art on the contrastive task, with accuracies slightly above 0.95 for both targeted phenomena. They also score highly on the four translation quality metrics, with COMET scores of 0.92 and 0.89. Moreover, we observed a “wise get wiser” effect, as the improvement brought by reasoning is positively correlated to the scores of the same LLMs without reasoning. In other words, the strongest models are also those that benefit the most from reasoning.
@@
8.  Acknowledgments
We are grateful to the Swiss National Science Foundation for its support through the EXOMAT grant n. 228494, External Knowledge for Low-resource Machine Translation. We acknowledge the support of HES-SO through the grant n. AGP-140146. We would like to warmly thank the anonymous LREC reviewers for their insightful suggestions.
@@ 9.  Bibliographical References
  • D. M. Alves, J. Pombal, N. M. Guerreiro, P. H. Martins, J. Alves, A. Farajian, B. Peters, R. Rei, P. Fernandes, S. Agrawal, P. Colombo, J. G. C. de Souza, and A. Martins (2024) Tower: an open multilingual large language model for translation-related tasks. In First Conference on Language Modeling, External Links: Link Cited by: §4.
  • R. Bawden, R. Sennrich, A. Birch, and B. Haddow (2018) Evaluating discourse phenomena in neural machine translation. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long Papers), New Orleans, Louisiana, pp. 1304–1313. External Links: Link, Document Cited by: §1, §2, §2, §3.1, Table 1, §5.2.
  • R. Bawden and F. Yvon (2023) Investigating the translation performance of a large multilingual language model: the case of BLOOM. In Proceedings of the 24th Annual Conference of the European Association for Machine Translation, Tampere, Finland, pp. 157–170. External Links: Link Cited by: §2.
  • R. Bawden (2018) Going beyond the sentence: contextual machine translation of dialogue. Ph.D. Thesis, Université Paris-Saclay. External Links: Link Cited by: §3.1.
  • E. Briakou, J. Luo, C. Cherry, and M. Freitag (2024) Translating step-by-step: decomposing the translation process forimproved translation quality of long-form texts. In Proceedings of the Ninth Conference on Machine Translation, Miami, Florida, USA, pp. 1301–1317. External Links: Link Cited by: footnote 8.
  • N. Campolungo, F. Martelli, F. Saina, and R. Navigli (2022) DiBiMT: a novel benchmark for measuring Word Sense Disambiguation biases in Machine Translation. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), S. Muresan, P. Nakov, and A. Villavicencio (Eds.), Dublin, Ireland, pp. 4331–4352. External Links: Link, Document Cited by: §2.
  • S. Castilho and R. Knowles (2025) A survey of context in neural machine translation and its evaluation. Natural Language Processing 31 (4), pp. 986–1016. External Links: Document, Link Cited by: §2.
  • S. Castilho, C. Q. Mallon, R. Meister, and S. Yue (2023) Do online machine translation systems care for context? What about a GPT model?. In Proceedings of the 24th Annual Conference of the European Association for Machine Translation, M. Nurminen, J. Brenner, M. Koponen, S. Latomaa, M. Mikhailov, F. Schierl, T. Ranasinghe, E. Vanmassenhove, S. A. Vidal, N. Aranberri, M. Nunziatini, C. P. Escartín, M. Forcada, M. Popovic, C. Scarton, and H. Moniz (Eds.), Tampere, Finland, pp. 393–417. External Links: Link Cited by: §2.
  • A. Chatterji, T. Cunningham, D. J. Deming, Z. Hitzig, C. Ong, C. Y. Shan, and K. Wadman (2025) How people use ChatGPT. NBER Working Paper 34255. External Links: Link Cited by: §2.
  • M. Cui, P. Gao, W. Liu, J. Luan, and B. Wang (2025) Multilingual machine translation with open large language models at practical scale: an empirical study. In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), L. Chiruzzo, A. Ritter, and L. Wang (Eds.), Albuquerque, New Mexico, pp. 5420–5443. External Links: Link, Document, ISBN 979-8-89176-189-6 Cited by: §1.
  • N. Dahan, R. Bawden, and F. Yvon (2024) Survey of automatic metrics for evaluating machine translation at the document level. MaTOS Deliverable D4-4.1. External Links: Link Cited by: §2.
  • L. Guillou, C. Hardmeier, E. Lapshinova-Koltunski, and S. Loáiciga (2018) A pronoun test suite evaluation of the English–German MT systems at WMT 2018. In Proceedings of the Third Conference on Machine Translation: Shared Task Papers, O. Bojar, R. Chatterjee, C. Federmann, M. Fishel, Y. Graham, B. Haddow, M. Huck, A. J. Yepes, P. Koehn, C. Monz, M. Negri, A. Névéol, M. Neves, M. Post, L. Specia, M. Turchi, and K. Verspoor (Eds.), Belgium, Brussels, pp. 570–577. External Links: Link, Document Cited by: §2.
  • L. Guillou and C. Hardmeier (2016) PROTEST: a test suite for evaluating pronouns in machine translation. In Proceedings of the Tenth International Conference on Language Resources and Evaluation (LREC’16), N. Calzolari, K. Choukri, T. Declerck, S. Goggi, M. Grobelnik, B. Maegaard, J. Mariani, H. Mazo, A. Moreno, J. Odijk, and S. Piperidis (Eds.), Portorož, Slovenia, pp. 636–643. External Links: Link Cited by: §2.
  • M. He, Y. Liu, S. Tao, Y. Luo, H. Zeng, C. Su, L. Zhang, H. Ma, D. Wei, W. Meng, et al. (2025) R1-T1: fully incentivizing translation capability in llms via reasoning learning. arXiv preprint arXiv:2502.19735. External Links: Link Cited by: §2.
  • A. Hendy, M. Abdelrehim, A. Sharaf, V. Raunak, M. Gabr, H. Matsushita, Y. J. Kim, M. Afify, and H. H. Awadalla (2023) How good are GPT models at machine translation? a comprehensive evaluation. arXiv preprint arXiv:2302.09210. External Links: Link Cited by: §2.
  • V. Iyer, P. Chen, and A. Birch (2023) Towards effective disambiguation for machine translation with large language models. In Proceedings of the Eighth Conference on Machine Translation, Singapore, pp. 482–495. External Links: Link Cited by: §2.
  • L. Jin, J. He, J. May, and X. Ma (2023) Challenges in context-aware neural machine translation. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, H. Bouamor, J. Pino, and K. Bali (Eds.), Singapore, pp. 15246–15263. External Links: Link, Document Cited by: §2.
  • M. Karpinska and M. Iyyer (2023) Large language models effectively leverage document-level context for literary translation, but critical errors persist. In Proceedings of the Eighth Conference on Machine Translation, Singapore, pp. 419–451. External Links: Link Cited by: §2.
  • T. Kocmi, E. Artemova, E. Avramidis, R. Bawden, O. Bojar, K. Dranch, A. Dvorkovich, S. Dukanov, M. Fishel, M. Freitag, T. Gowda, R. Grundkiewicz, B. Haddow, M. Karpinska, P. Koehn, H. Lakougna, J. Lundin, C. Monz, K. Murray, M. Nagata, S. Perrella, L. Proietti, M. Popel, M. Popović, P. Riley, M. Shmatova, S. Steingrímsson, L. Yankovskaya, and V. Zouhar (2025) Findings of the WMT25 general machine translation shared task: time to stop evaluating on easy test sets. In Proceedings of the Tenth Conference on Machine Translation (WMT 2025), Suzhou, China, pp. 355–413. External Links: Link Cited by: §2.
  • T. Kocmi, E. Avramidis, R. Bawden, O. Bojar, A. Dvorkovich, C. Federmann, M. Fishel, M. Freitag, T. Gowda, R. Grundkiewicz, B. Haddow, M. Karpinska, P. Koehn, B. Marie, C. Monz, K. Murray, M. Nagata, M. Popel, M. Popović, M. Shmatova, S. Steingrímsson, and V. Zouhar (2024) Findings of the WMT24 general machine translation shared task: the LLM era is here but MT is not solved yet. In Proceedings of the Ninth Conference on Machine Translation, B. Haddow, T. Kocmi, P. Koehn, and C. Monz (Eds.), Miami, Florida, USA, pp. 1–46. External Links: Link, Document Cited by: §1, §2.
  • T. Kocmi, E. Avramidis, R. Bawden, O. Bojar, A. Dvorkovich, C. Federmann, M. Fishel, M. Freitag, T. Gowda, R. Grundkiewicz, B. Haddow, P. Koehn, B. Marie, C. Monz, M. Morishita, K. Murray, M. Nagata, T. Nakazawa, M. Popel, M. Popović, and M. Shmatova (2023) Findings of the 2023 conference on machine translation (WMT23): LLMs are here but not quite there yet. In Proceedings of the Eighth Conference on Machine Translation, Singapore, pp. 1–42. External Links: Link Cited by: §2.
  • T. Kocmi and C. Federmann (2023) Large language models are state-of-the-art evaluators of translation quality. arXiv preprint arXiv:2302.14520. External Links: Link Cited by: §2.
  • D. Larionov, S. Takeshita, R. Zhang, Y. Chen, C. Leiter, Z. Wang, C. Greisinger, and S. Eger (2025) DeepSeek vs. o3-mini: how well can reasoning LLMs evaluate MT and summarization?. arXiv preprint arXiv:2504.08120. External Links: Link Cited by: §2.
  • S. Läubli, R. Sennrich, and M. Volk (2018) Has machine translation achieved human parity? A case for document-level evaluation. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing, E. Riloff, D. Chiang, J. Hockenmaier, and J. Tsujii (Eds.), Brussels, Belgium, pp. 4791–4796. External Links: Link, Document Cited by: §2.
  • S. Liu, C. Lyu, M. Wu, L. Wang, W. Luo, K. Zhang, and Z. Shang (2025) New trends for modern machine translation with large reasoning models. arXiv preprint arXiv:2503.10351. External Links: Link Cited by: §2.
  • A. Lopes, M. A. Farajian, R. Bawden, M. Zhang, and A. F. T. Martins (2020) Document-level neural MT: a systematic comparison. In Proceedings of the 22nd Annual Conference of the European Association for Machine Translation, A. Martins, H. Moniz, S. Fumega, B. Martins, F. Batista, L. Coheur, C. Parra, I. Trancoso, M. Turchi, A. Bisazza, J. Moorkens, A. Guerberof, M. Nurminen, L. Marg, and M. L. Forcada (Eds.), Lisboa, Portugal, pp. 225–234. External Links: Link Cited by: §2, §5.2.
  • S. Maruf, F. Saleh, and G. Haffari (2021) A survey on document-level neural machine translation: methods and evaluation. ACM Computing Surveys 54 (2). External Links: ISSN 0360-0300, Link, Document Cited by: §2.
  • L. Miculicich Werlen and A. Popescu-Belis (2017) Validation of an automatic metric for the accuracy of pronoun translation (APT). In Proceedings of the Third Workshop on Discourse in Machine Translation, B. Webber, A. Popescu-Belis, and J. Tiedemann (Eds.), Copenhagen, Denmark, pp. 17–25. External Links: Link, Document Cited by: §2, §3.2.
  • M. Müller, A. Rios, E. Voita, and R. Sennrich (2018) A large-scale test set for the evaluation of context-aware pronoun translation in neural machine translation. In Proceedings of the Third Conference on Machine Translation: Research Papers, O. Bojar, R. Chatterjee, C. Federmann, M. Fishel, Y. Graham, B. Haddow, M. Huck, A. J. Yepes, P. Koehn, C. Monz, M. Negri, A. Névéol, M. Neves, M. Post, L. Specia, M. Turchi, and K. Verspoor (Eds.), Brussels, Belgium, pp. 61–72. External Links: Link, Document Cited by: §2.
  • NLLB-Team, M. R. Costa-jussà, J. Cross, O. Çelebi, M. Elbayad, K. Heafield, K. Heffernan, E. Kalbassi, J. Lam, D. Licht, J. Maillard, A. Sun, S. Wang, G. Wenzek, A. Youngblood, B. Akula, L. Barrault, G. M. Gonzalez, P. Hansanti, J. Hoffman, S. Jarrett, K. R. Sadagopan, D. Rowe, S. Spruit, C. Tran, P. Andrews, N. F. Ayan, S. Bhosale, S. Edunov, A. Fan, C. Gao, V. Goswami, F. Guzmán, P. Koehn, A. Mourachko, C. Ropers, S. Saleem, H. Schwenk, and J. Wang (2022) No language left behind: scaling human-centered machine translation. arXiv 2207.04672. External Links: Link Cited by: §4.
  • P. Pal, A. Birch, and K. Heafield (2024) Document-level machine translation with large-scale public parallel corpora. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), L. Ku, A. Martins, and V. Srikumar (Eds.), Bangkok, Thailand, pp. 13185–13197. External Links: Link, Document Cited by: §5.2.
  • K. Papineni, S. Roukos, T. Ward, and W. Zhu (2002) BLEU: a method for automatic evaluation of machine translation. In Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics, P. Isabelle, E. Charniak, and D. Lin (Eds.), Philadelphia, Pennsylvania, USA, pp. 311–318. External Links: Link, Document Cited by: §2.
  • J. Pilault, X. Garcia, A. Bražinskas, and O. Firat (2023) Interactive-chain-prompting: ambiguity resolution for crosslingual conditional generation with interaction. In Proceedings of the 13th International Joint Conference on Natural Language Processing and the 3rd Conference of the Asia-Pacific Chapter of the Association for Computational Linguistics (Volume 1: Long Papers), J. C. Park, Y. Arase, B. Hu, W. Lu, D. Wijaya, A. Purwarianti, and A. A. Krisnadhi (Eds.), Nusa Dua, Bali, pp. 455–483. External Links: Link, Document Cited by: §2.
  • A. Popescu-Belis (2019) Context in neural machine translation: a review of models and evaluations. arXiv preprint arXiv:1901.09115. External Links: Link Cited by: §2.
  • M. Popović (2015) ChrF: character n-gram F-score for automatic MT evaluation. In Proceedings of the Tenth Workshop on Statistical Machine Translation, O. Bojar, R. Chatterjee, C. Federmann, B. Haddow, C. Hokamp, M. Huck, V. Logacheva, and P. Pecina (Eds.), Lisbon, Portugal, pp. 392–395. External Links: Link, Document Cited by: §2.
  • M. Post and M. Junczys-Dowmunt (2024) Evaluation and large-scale training for contextual machine translation. In Proceedings of the Ninth Conference on Machine Translation, Miami, Florida, USA, pp. 1125–1139. External Links: Link Cited by: §6.
  • R. Rei, C. Stewart, A. Farinha, and A. Lavie (2020) COMET: a neural framework for MT evaluation. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), Online, pp. 2685–2702. External Links: Link, Document Cited by: §2.
  • A. Rios, M. Müller, and R. Sennrich (2018) The word sense disambiguation test suite at WMT18. In Proceedings of the Third Conference on Machine Translation: Shared Task Papers, O. Bojar, R. Chatterjee, C. Federmann, M. Fishel, Y. Graham, B. Haddow, M. Huck, A. J. Yepes, P. Koehn, C. Monz, M. Negri, A. Névéol, M. Neves, M. Post, L. Specia, M. Turchi, and K. Verspoor (Eds.), Belgium, Brussels, pp. 588–596. External Links: Link, Document Cited by: §2.
  • A. Rios Gonzales, L. Mascarell, and R. Sennrich (2017) Improving word sense disambiguation in neural machine translation with sense embeddings. In Proceedings of the Second Conference on Machine Translation, O. Bojar, C. Buck, R. Chatterjee, C. Federmann, Y. Graham, B. Haddow, M. Huck, A. J. Yepes, P. Koehn, and J. Kreutzer (Eds.), Copenhagen, Denmark, pp. 11–19. External Links: Link, Document Cited by: §2.
  • J. Tiedemann and Y. Scherrer (2017) Neural machine translation with extended context. In Proceedings of the Third Workshop on Discourse in Machine Translation, B. Webber, A. Popescu-Belis, and J. Tiedemann (Eds.), Copenhagen, Denmark, pp. 82–92. External Links: Link, Document Cited by: §5.2.
  • G. Vernikos, B. Thompson, P. Mathur, and M. Federico (2022) Embarrassingly easy document-level MT metrics: how to convert any pretrained metric into a document-level metric. In Proceedings of the Seventh Conference on Machine Translation (WMT), P. Koehn, L. Barrault, O. Bojar, F. Bougares, R. Chatterjee, M. R. Costa-jussà, C. Federmann, M. Fishel, A. Fraser, M. Freitag, Y. Graham, R. Grundkiewicz, P. Guzman, B. Haddow, M. Huck, A. Jimeno Yepes, T. Kocmi, A. Martins, M. Morishita, C. Monz, M. Nagata, T. Nakazawa, M. Negri, A. Névéol, M. Neves, M. Popel, M. Turchi, and M. Zampieri (Eds.), Abu Dhabi, United Arab Emirates (Hybrid), pp. 118–128. External Links: Link Cited by: §5.2.
  • D. Vilar, M. Freitag, C. Cherry, J. Luo, V. Ratnakar, and G. Foster (2023) Prompting PaLM for translation: assessing strategies and performance. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), A. Rogers, J. Boyd-Graber, and N. Okazaki (Eds.), Toronto, Canada, pp. 15406–15427. External Links: Link, Document Cited by: §2.
  • E. Voita, R. Sennrich, and I. Titov (2019) When a good translation is wrong in context: context-aware machine translation improves on deixis, ellipsis, and lexical cohesion. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, A. Korhonen, D. Traum, and L. Màrquez (Eds.), Florence, Italy, pp. 1198–1212. External Links: Link, Document Cited by: §2.
  • L. Wang, C. Lyu, T. Ji, Z. Zhang, D. Yu, S. Shi, and Z. Tu (2023) Document-level machine translation with large language models. arXiv preprint arXiv:2304.02210. External Links: Link Cited by: §2.
  • J. Wei, X. Wang, D. Schuurmans, M. Bosma, F. Xia, E. Chi, Q. V. Le, D. Zhou, et al. (2022) Chain-of-thought prompting elicits reasoning in large language models. Advances in Neural Information Processing Systems (NeurIPS) 35, pp. 24824–24837. External Links: Link Cited by: §1, §2.
  • M. Wu, T. Vu, L. Qu, G. Foster, and G. Haffari (2024) Adapting large language models for document-level machine translation. arXiv preprint arXiv:2401.06468. External Links: Link Cited by: §2.
  • Y. Ye, B. Fu, C. Huang, Y. Chen, and X. Shi (2025) How well do large reasoning models translate? A comprehensive evaluation for multi-domain machine translation. arXiv preprint arXiv:2505.19987. External Links: Link Cited by: §2.
  • B. Zhang, A. Bapna, M. Johnson, A. Dabirmoghaddam, N. Arivazhagan, and O. Firat (2022) Multilingual document-level translation enables zero-shot transfer from sentences to documents. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), S. Muresan, P. Nakov, and A. Villavicencio (Eds.), Dublin, Ireland, pp. 4176–4192. External Links: Link, Document Cited by: §5.2.
  • B. Zhang, B. Haddow, and A. Birch (2023) Prompting large language model for machine translation: a case study. In Proceedings of the 40th International Conference on Machine Learning, A. Krause, E. Brunskill, K. Cho, B. Engelhardt, S. Sabato, and J. Scarlett (Eds.), Proceedings of Machine Learning Research, Vol. 202, pp. 41092–41110. External Links: Link Cited by: §2.
  • T. Zhang, V. Kishore, F. Wu, K. Q. Weinberger, and Y. Artzi (2020) BERTScore: evaluating text generation with BERT. In Proceedings of the International Conference on Learning Representations (ICLR), External Links: Link Cited by: §2.
@@ A.1.  Prompt Variants
This section presents the full text of the prompts used in our contrastive evaluation. Following common prompt engineering practice, we have experimented with various system prompts (or none) and user prompts. The various prompts differ in their structures and instruction levels, while maintaining a consistent task format. Placeholders appear in curly braces in the prompt templates: {context_en}, {source_sentence}, {context_fr}, {option1}, {option2}. In the contrastive setting, two candidate translations are presented, and the model must output a constrained choice, either as “Choice: 1” or as “Choice: 2”.
No system prompt, simple user prompt
@@
Simple system and user prompts
@@
Detailed system prompt, simple user prompt
@@
Simple system prompt, step-by-step (reasoning) user prompt for anaphora
@@
Simple system prompt, step-by-step (reasoning) user prompt for lexical choice
@@ A.2.  Results: Anaphora
Below are the results on the contrastive anaphora benchmark for each LLM and prompt type, in terms of accuracy and inconsistency scores, elapsed time and cost of LLM.
System and user prompts
Model None & simple Both simple Detailed & simple) Simple & step-by-step
@@
Table 7: Accuracy ($\uparrow$) and inconsistency ($\downarrow$) across models and prompts on the anaphora validation set (separated by ‘/’ in each cell). The systems are ranked by decreasing accuracy of the best overall prompt, which is the fourth one.
System and user prompts
Model None & simple Both simple Detailed & simple) Simple & step-by-step
@@
Table 8: Mean elapsed time per prompt in seconds, across prompt configurations on the anaphora validation set. Open-source (Ollama) models listed first, followed by OpenAI models.
System and user prompts
Model None & simple Both simple Detailed & simple) Simple & step-by-step
@@
Table 9: Total API cost (USD) by prompt configuration for OpenAI models on the anaphora validation set. Experiments were run in summer 2025; reported costs reflect that period’s pricing and our token usage.
@@
Table 10: Comparison of accuracy (Acc. $\uparrow$) and inconsistency (Inc. $\downarrow$) on validation and test sets for the anaphora task, across prompt configurations. Validation uses the first 100 couples and reproduces the scores from Table 7 above; test uses the held-out 100 couples. Each couple is presented in both option orders (correct option first/second), and scores are averaged across the two presentations.